What is the potential in volts at a distance of 140 pm from a Ti4+ ion? (The answer should be written as a number with two digits to the right of the decimal and with the correct sign.)
Answer #1

Solution

Answer: +41.14 Volts

Assuming the Ti4+ ion to be a point charge of charge ,

Q=+4e=4*1.6*10^{-19}C=6.4*10^{-19} CElectronic charge is e=1.6*10^{-19} C

Thus potential at a distance of,

R=140pm=140*10^{-12}m

Is,

V=\frac{1}{4\pi \epsilon_0}\frac{Q}{R}=(9*10^9)*\frac{6.4*10^{-19}}{140*10^{-12}}=41.143Volts(Standard value of 1/4pi epsilon is 9*10^9 m/F)

Thus rounding off to two decimal place,

V=41.14Volts